# DyPCL: Dynamic Phoneme-level Contrastive Learning for Dysarthric Speech Recognition

### v1

@@ Abstract
Dysarthric speech recognition often suffers from performance degradation due to the intrinsic diversity of dysarthric severity and extrinsic disparity from normal speech. To bridge these gaps, we propose a Dynamic Phoneme-level Contrastive Learning (DyPCL) method, which leads to obtaining invariant representations across diverse speakers. We decompose the speech utterance into phoneme segments for phoneme-level contrastive learning, leveraging dynamic connectionist temporal classification alignment. Unlike prior studies focusing on utterance-level embeddings, our granular learning allows discrimination of subtle parts of speech. In addition, we introduce dynamic curriculum learning, which progressively transitions from easy negative samples to difficult-to-distinguishable negative samples based on phonetic similarity of phoneme. Our approach to training by difficulty levels alleviates the inherent variability of speakers, better identifying challenging speeches. Evaluated on the UASpeech dataset, DyPCL outperforms baseline models, achieving an average 22.10% relative reduction in word error rate (WER) across the overall dysarthria group.
@@ 3.1 Phoneme-level Contrastive Learning
This process yields roughly 48.65 billion triplet pairs. We use stratified sampling to balance training cost and efficiency, limiting each anchor to a maximum of five positive samples, and each anchor-positive pair to five negative samples. Consequently, 1.18 million triplet pairs are created for contrastive learning. During training, we randomly sampled 200,000 triplet pairs.
@@ 3.3 Dynamic Curriculum Learning
In the DyPCL framework, we dynamically select negative samples using a curriculum learning approach. This dynamic selection strategy gradually increases the difficulty of negative samples, fostering a more robust learning process. The difficulty is determined by phoneme distance, which is measured using the PanPhon tool Mortensen et al. (2016)  [^2] https://github.com/dmort27/panphon. The hamming feature edit distance is used to calculate phoneme distance, with equal weighting across all 21 articulatory features that define each phoneme. The phoneme distance reflects how similar or different phonemes sound phonetically, ranging from 0.0416 (most similar) to 0.583 (most different). Figure 6 in Appendix A shows phoneme distance matrix in heat map. By learning to distinguish similar-sounding phonemes in the embedding space through DyPCL, the model can further improve recognition accuracy for given phonemes.
@@ 5.5 Comparison with Benchmarks
Table 4 presents a comparison of the performance of our model against state-of-the-art (SOTA) methods on the CTEST set. Bhat et al. (2022a) employs a two-stage augmentation approach. The second and third methods come from Wang et al. (2024b), which was the first to introduce contrastive learning for UASpeech. In their study, the "Speaker Dependent (SD) w/finetune" method focuses on fine-tuning speaker-specific word prototypes, while the "Speaker Independent (SI)" method works across both word- and speaker-level instances to improve generalization. The SD approach improved the WER to 13. 49%, while the SI method further reduced it to 12. 09%. In contrast, our model, DyPCL (GP), achieved the lowest WER at severity levels, significantly reducing the overall WER to 10. 10. 34%, outperforming all previous models.
Table 4 also compares our model, DyPCL (GP), against other methods on the TEST set. For a fair comparison, the results from previous works are reported without data augmentation (DA), focusing on the core contributions of each method. Wang et al. (2023b) used hyperparameter adaptation to handle speaker differences, achieving an overall WER of 30.49%. Hsieh and Wu (2024a) applied curriculum learning, training progressively from high to low intelligibility groups with their propose re-grouping method, but still reported a relatively high WER of 7.99% for the easiest group (H), highlighting challenges in early-stage learning. Both Hu et al. (2023) and Hu et al. (2024) used speaker-adaptive training, incorporating speaker-specific articulatory and acoustic features, achieving WERs of 24.62% and 22.62%, respectively. Geng et al. (2023a) integrated severity information and system combination, and without DA, reported an overall WER of 24.57%.
Our model, DyPCL (GP), not only achieved the lowest overall WER of 20.7%, outperforming all previous methods, but also demonstrated significant improvements for the low intelligibility groups. With WERs of 17.6% for the L group and 49.45% for the VL group, it maintained strong performance across all dysarthria severity levels, achieving a WER of just 2.77% for the High (H) intelligibility group. This level of robustness highlights the reliability of its performance.
@@ 6 Conclusion
This paper introduced the Dynamic Phoneme-level Contrastive Learning (DyPCL) framework to improve dysarthric speech recognition. DyPCL effectively tackles phoneme alignment challenges and accounting for phonetic difficulty through dynamic CTC alignment and curriculum learning. Our experiments on the UASpeech dataset demonstrated the effectiveness of DyPCL, which reduced the WER from 58.49% to 49.45% in the Very Low (VL) intelligibility group, and the overall WER across all dysarthria group from 25.97% to 20.23%. These results underscore DyPCL’s capability to capture subtle phonetic variations, significantly enhancing speech recognition accuracy across all levels of dysarthria severity

### v2

@@ Abstract
Dysarthric speech recognition often suffers from performance degradation due to the intrinsic diversity of dysarthric severity and extrinsic disparity from normal speech. To bridge these gaps, we propose a Dynamic Phoneme-level Contrastive Learning (DyPCL) method, which leads to obtaining invariant representations across diverse speakers. We decompose the speech utterance into phoneme segments for phoneme-level contrastive learning, leveraging dynamic connectionist temporal classification alignment. Unlike prior studies focusing on utterance-level embeddings, our granular learning allows discrimination of subtle parts of speech. In addition, we introduce dynamic curriculum learning, which progressively transitions from easy negative samples to difficult-to-distinguish negative samples based on phonetic similarity of the phoneme. Our approach to training by difficulty levels alleviates the inherent variability of speakers, better identifying challenging speeches. Evaluated on the UASpeech dataset, DyPCL outperforms baseline models, achieving an average 22.10% relative reduction in word error rate (WER) across the overall dysarthria group.
@@ 3.1 Phoneme-level Contrastive Learning
This process yields roughly 48.65 billion triplet pairs. We use stratified sampling to balance training cost and efficiency, limiting each anchor to a maximum of five positive samples and each anchor-positive pair to five negative samples. Consequently, 1.18 million triplet pairs are created for contrastive learning. During training, we randomly sampled 200,000 triplet pairs.
@@ 3.3 Dynamic Curriculum Learning
In the DyPCL framework, we dynamically select negative samples using a curriculum learning approach. This dynamic selection strategy gradually increases the difficulty of negative samples, fostering a more robust learning process. The difficulty is determined by phoneme distance, which is measured using the PanPhon tool Mortensen et al. (2016)  [^2] https://github.com/dmort27/panphon. The hamming feature edit distance is used to calculate phoneme distance, with equal weighting across all 21 articulatory features that define each phoneme. The phoneme distance reflects how similar or different phonemes sound phonetically, ranging from 0.0416 (most similar) to 0.583 (most different). Figure 6 in Appendix A shows the phoneme distance matrix in a heat map. By learning to distinguish similar-sounding phonemes in the embedding space through DyPCL, the model can further improve recognition accuracy for given phonemes.
@@ 5.5 Comparison with Benchmarks
Table 4 presents a comparison of the performance of our model against state-of-the-art (SOTA) methods on the CTEST set. Bhat et al. (2022a) employs a two-stage augmentation approach. The second and third methods come from Wang et al. (2024b), which was the first to introduce contrastive learning for UASpeech. In their study, the "Speaker Dependent (SD) w/ finetune" method focuses on fine-tuning speaker-specific word prototypes, while the "Speaker Independent (SI)" method works across both word- and speaker-level instances to improve generalization. The SD approach improved the WER to 13.49%, while the SI method further reduced it to 12.09%. In contrast, our model, DyPCL (GP), achieved the lowest WER at severity levels, significantly reducing the overall WER to 10.34%, outperforming all previous models.
Table 4 also compares our model, DyPCL (GP), against other methods on the TEST set. For a fair comparison, the results from previous works are reported without data augmentation (DA), focusing on the core contributions of each method. Wang et al. (2023b) used hyperparameter adaptation to handle speaker differences, achieving an overall WER of 30.49%. Hsieh and Wu (2024a) applied curriculum learning, training progressively from high to low intelligibility groups with their proposed re-grouping method, but still reported a relatively high WER of 7.99% for the easiest group (H), highlighting challenges in early-stage learning. Both Hu et al. (2023) and Hu et al. (2024) used speaker-adaptive training, incorporating speaker-specific articulatory and acoustic features, achieving WERs of 24.62% and 22.62%, respectively. Geng et al. (2023a) integrated severity information and system combination, and without DA, reported an overall WER of 24.57%.
Our model, DyPCL (GP), not only achieved the lowest overall WER of 20.7%, outperforming all previous methods but also demonstrated significant improvements for the low intelligibility groups. With WERs of 17.6% for the L group and 49.45% for the VL group, it maintained strong performance across all dysarthria severity levels, achieving a WER of just 2.77% for the High (H) intelligibility group. This level of robustness highlights the reliability of its performance.
@@ 6 Conclusion
This paper introduced the Dynamic Phoneme-level Contrastive Learning (DyPCL) framework to improve dysarthric speech recognition. DyPCL effectively tackles phoneme alignment challenges and accounts for phonetic difficulty through dynamic CTC alignment and curriculum learning. Our experiments on the UASpeech dataset demonstrated the effectiveness of DyPCL, which reduced the WER from 58.49% to 49.45% in the Very Low (VL) intelligibility group and the overall WER across all dysarthria groups from 25.97% to 20.23%. These results underscore DyPCL’s capability to capture subtle phonetic variations, significantly enhancing speech recognition accuracy across all levels of dysarthria severity